# SynQuE: Estimating Synthetic Dataset Quality Without Annotations

### v1

@@ Abstract
We introduce and formalize the Synthetic Dataset Quality Estimation (SynQuE) problem: ranking synthetic datasets by their expected real-world task performance using only limited unannotated real data. This addresses a critical and open challenge where data is scarce due to collection costs or privacy constraints. We establish the first comprehensive benchmarks for this problem by introducing and evaluating proxy metrics that choose synthetic data for training to maximize task performance on real data. We introduce the first proxy metrics for SynQuE by adapting distribution and diversity-based distance measures to our context via embedding models. To address the shortcomings of these metrics on complex planning tasks, we propose Lens, a novel proxy that leverages large language model reasoning. Our results show that SynQuE proxies correlate with real task performance across diverse tasks, including sentiment analysis, Text2SQL, web navigation, and image classification, with Lens consistently outperforming others on complex tasks by capturing nuanced characteristics. For instance, on text-to-SQL parsing, training on the top-3 synthetic datasets selected via SynQuE proxies can raise accuracy from 30.4% to 38.4 (+8.1)% on average compared to selecting data indiscriminately. This work establishes SynQuE as a practical framework for synthetic data selection under real-data scarcity and motivates future research on foundation model-based data characterization and fine-grained data selection.
@@ 1 Introduction
Figure 1: SynQuE uses synthetic data and unlabeled samples of real data to estimate synthetic data quality. Proxy scores are used to rank and select the datasets that lead to the best task performance
@@ 2 Related Work
Evaluating data quality
Metrics like Proxy-$\mathcal{A}$-Distance ($\mathbf{PAD}$) estimate domain divergence by training classifiers to distinguish source and target data (ben-david_theory_2010; ben-david_analysis_2006; quinonero-candela_dataset_2022). While effective in some NLP tasks (elsahar_annotate_2019), $\mathbf{PAD}$ struggles in noisy, complex settings such as agent planning (he_webvoyager_2024). Diversity metrics like mean-distance-to-medoids ($\mathbf{MDM}$)(cox_directed_2021) and DCScore(zhu_measuring_2025) approximate the coverage of synthetic data. However, these measures do not capture common synthetic data factual inconsistencies from generative model hallucinations (huang_survey_2025; li_evaluating_2023; gunjal_detecting_2023), which can harm downstream performance (geng_unmet_2025; yu_large_2023; casco-rodriguez_self-consuming_2023; li_synthetic_2023; hataya_will_2023; briesch_large_2023). Other works focus on data selection using scaling laws by training a regression model on the results of short training runs to predict the best data mixture on a labeled validation set (liu_regmix_2025; magnusson_datadecide_2025). This approach, while effective, is computationally expensive and fundamentally requires labeled real data, which is unavailable in the SynQuE setting. Finally, another group of works try to distinguish between human and machine text using divergence measures (pillutla_mauve_2021) or LLM-as-a-judge (gu_survey_2025; krumdick_no_2025; zheng_judging_2023). Neither of these two techniques have been studied for synthetic data selection. Furthermore, the latter also suffers from limitations of generative models previously mentioned such as hallucination. Our work establishes the first systematic benchmark and results for synthetic data selection in a practical, fully unsupervised regime.
@@ 5.1 Results Analysis
Computers .22 .22 .36 .65 .28 .80 -.39 -.65 .19 .65 .64 .38 .51 .78 .34 .87 -.39 -.62 .87 .78
Apps .29 .48 -.20 -.23 .40 .77 -.44 -.56 .65 .70 .64 .66 .68 .76 .57 .80 -.44 -.56 .32 .65
Movies .33 .48 .33 .36 .52 .57 -.60 -.40 -.14 -.35 .43 .49 .57 .37 .81 .56 -.67 -.45 .81 .60
Average .28 .39 .16 .26 .40 .71 -.47 -.54 .23 .33 .57 .51 .59 .64 .57 .74 -.50 -.55 .67 .68
@@ B.1 Ablation study on Lens preferential bias
B.2 Ablation study on Lens rubric design
@@
B.3 Ablation study on MMD2 kernel functions
Table 10: We report Pearson / Spearman correlation coefficients. The Laplacian kernel performs best on text tasks, but all kernels show limited effectiveness on more complex domains.

### v3

@@ Abstract
We introduce and formalize the Synthetic Dataset Quality Estimation (SynQuE) problem: ranking synthetic datasets by their expected real-world task performance using only limited unannotated real data. This addresses a critical and open challenge where data is scarce due to collection costs or privacy constraints. We establish the first comprehensive benchmarks for this problem by introducing and evaluating proxy metrics that choose synthetic data for training to maximize task performance on real data. We introduce the first proxy metrics for SynQuE by adapting distribution and diversity-based distance measures to our context via embedding models. To address the shortcomings of these metrics on complex planning tasks, we propose Lens, a novel proxy that leverages large language model reasoning. Our results show that SynQuE proxies correlate with real task performance across diverse tasks, including sentiment analysis, Text2SQL, web navigation, and image classification, with Lens consistently outperforming others on complex tasks by capturing nuanced characteristics. For instance, on text-to-SQL parsing, training on the top-3 synthetic datasets selected via SynQuE proxies can raise accuracy from 30.4% to 38.4 (+8.1)% on average compared to selecting data indiscriminately. This work establishes SynQuE as a practical framework for synthetic data selection under real-data scarcity and motivates future research on foundation model-based data characterization and fine-grained data selection. We release our code [^1] https://github.com/r2llab/SynQuE.
@@ 1 Introduction
Figure 1: SynQuE uses synthetic data and unlabeled samples of real data to estimate synthetic data quality. Proxy scores are used to rank and select the datasets that lead to the best task performance.
@@ 2 Related Work
Distributional and diversity metrics
Metrics like Proxy-$\mathcal{A}$-Distance ($\mathbf{PAD}$) estimate domain divergence by training classifiers to distinguish source and target data (ben-david_theory_2010; ben-david_analysis_2006; quinonero-candela_dataset_2022). While effective in some NLP tasks (elsahar_annotate_2019), $\mathbf{PAD}$ struggles in noisy, complex settings such as agent planning (he_webvoyager_2024). Diversity metrics like mean-distance-to-medoids ($\mathbf{MDM}$)(cox_directed_2021) and DCScore(zhu_measuring_2025) approximate the coverage of synthetic data. However, these measures do not capture common synthetic data factual inconsistencies from generative model hallucinations (huang_survey_2025; li_evaluating_2023; gunjal_detecting_2023), which can harm downstream performance (geng_unmet_2025; yu_large_2023; casco-rodriguez_self-consuming_2023; li_synthetic_2023; hataya_will_2023; briesch_large_2023).
Data selection and weighting
A growing body of work focuses on selecting high-value training samples. Some approaches utilize scaling laws by training regression models on short runs to predict the best data mixture (liu_regmix_2025; magnusson_datadecide_2025). Others employ influence functions or gradient-based heuristics to select influential subsets (zhang_harnessing_2024; hu_most_2025), or utilize small real datasets to learn instance-level weights for synthetic data (kuo_not_2025). Crucially, these methods focus on instance-level selection or weighting, often requiring access to labeled validation sets to compute gradients or train selection models. In contrast, SynQuE addresses dataset-level ranking in a zero-shot, gradient-free manner, making it suitable for “cold-start” settings where no labeled data exists. Finally, another group of works try to distinguish between human and machine text using divergence measures (pillutla_mauve_2021) or LLM-as-a-judge (gu_survey_2025; krumdick_no_2025; zheng_judging_2023). Neither of these two techniques have been studied for synthetic data selection. Furthermore, the latter also suffers from limitations of generative models previously mentioned such as hallucination. Our work establishes the first systematic benchmark and results for synthetic data selection in a practical, fully unsupervised regime.
@@ 5.1 Results Analysis
Computers .22 .22 .36 .65 .28 .80 .39 .65 .19 .65 .64 .38 .51 .78 .34 .87 .39 .62 .87 .78
Apps .29 .48 -.20 -.23 .40 .77 .44 .56 .65 .70 .64 .66 .68 .76 .57 .80 .44 .56 .32 .65
Movies .33 .48 .33 .36 .52 .57 .60 .40 -.14 -.35 .43 .49 .57 .37 .81 .56 .67 .45 .81 .60
Average .28 .39 .16 .26 .40 .71 .47 .54 .23 .33 .57 .51 .59 .64 .57 .74 .50 .55 .67 .68
@@ B.1 Ablation study on Lens preferential bias
B.2 Lens with different backbone models and sizes
Table 8: Comparison of original and debiased Lens scores for ranking synthetic datasets across various backbone models and sizes. For GPT-4.1, sentiment analysis results are reported only due to cost constraints.
LENS (Qwen2.5-7B-Instruct) Debiased Original
Task Spearman Pearson Spearman Pearson
Sentiment .25 .33 .26 .17
Text2SQL .33 .33 .31 .28
LENS (Qwen2.5-32B-Instruct) Debiased Original
Task Spearman Pearson Spearman Pearson
Sentiment .38 .26 .24 .23
Text2SQL .46 .43 .43 .44
LENS (Ministral-8B-Instruct) Debiased Original
Task Spearman Pearson Spearman Pearson
Sentiment .28 .39 .52 .31
Text2SQL .32 .29 .16 .16
LENS (Granite-8B-Instruct) Debiased Original
Task Spearman Pearson Spearman Pearson
Sentiment .21 .30 -.04 -.02
Text2SQL .22 .20 .19 .23
LENS (Granite-4.0-h-small) (32B) Debiased Original
Task Spearman Pearson Spearman Pearson
Sentiment .16 .39 .50 .44
Text2SQL .32 .36 .40 .48
LENS (Gemma3-12B-Instruct) Debiased Original
Task Spearman Pearson Spearman Pearson
Sentiment .10 .13 .30 .18
Text2SQL .30 .34 .25 .25
LENS (Gemma3-27B-Instruct) Debiased Original
Task Spearman Pearson Spearman Pearson
Sentiment .17 .33 .37 .31
Text2SQL .31 .38 .31 .27
LENS (GPT-4.1-2025-04-14) Debiased Original
Task Spearman Pearson Spearman Pearson
Sentiment .07 .30 .28 .37
As shown in table 8, Lens demonstrates robust ranking capabilities across a diverse range of backbone models and sizes. However, the impact of the debiasing step varies significantly by task complexity and model capability:
  • Text2SQL (Complex Generation): Debiasing consistently improves or maintains correlation strength across nearly all models. For instance, with Ministral-8B, debiasing doubles the Spearman correlation (.16 $\rightarrow$ .32), and it yields the highest overall performance with Qwen2.5-32B (.46). This suggests that for complex structural tasks, removing distributional priors remains beneficial even for capable models.
  • Sentiment Analysis (Simple Classification): The results are more nuanced. For models with severe initial bias, such as Granite-8B-Instruct, debiasing is critical, correcting a negative correlation (-.04) to a positive one (.21). However, for highly capable or larger models (e.g., GPT-4.1, Granite-4.0 (32B), Gemma3-27B), debiasing tends to degrade performance. We hypothesize that these stronger models are already well-calibrated for simple classification tasks; consequently, our analytical debiasing term may introduce unnecessary noise that overrides the model’s accurate intrinsic priors.
B.3 Ablation study on Lens rubric design
@@
Our ablation results from table 9 demonstrate the effectiveness of the rubric component of Lens. Table 10 demonstrates that generating too few or too many points will degrade the performance of Lens.
B.4 Ablation study on $\mathbf{MMD}^{2}$ kernel functions
Table 11: Pearson / Spearman correlation coefficients of $\mathbf{MMD}^{2}$ with different kernel functions.
@@ B.4 Ablation study on $\mathbf{MMD}^{2}$ kernel functions
Based on the ablation results from table 11, the Laplacian kernel performs best on text tasks, but all kernels show limited effectiveness on more complex domains.
B.5 Ablation study on $\mathbf{MDM}$ with different number of clusters
Table 12: Pearson / Spearman coefficients of $\mathbf{MDM}$ with different number of clusters (K=3, 5, 10, 20) on sentiment analysis.
K=3 K=5 K=10 K=20
.85/ .68 .86/ .68 .86/ .69 .86/ .69
The table with varying number of medoid clusters for $\mathbf{MDM}$ shows that changing the number of clusters has negligible effect on the correlation between $\mathbf{MDM}$ and F1 score.
B.6 Ablation study on using different encoder for representation-based metrics
Table 13: Pearson / Spearman correlation coefficients of proxies on sentiment analysis with BGE-M3 as the encoder.
Mauve $\mathbf{PAD}$ $\mathbf{MMD}^{2}$ $\mathbf{MDM}$
.60/ .65 .75/ .65 .82/ .67 .86/ .74
The results with BGE-M3777https://huggingface.co/BAAI/bge-m3 demonstrate that representation-based metrics are measuring fundamental distribution shifts rather than artifacts of a specific encoder.
B.7 Correlation of $\mathbf{PAD}$ predictions from different classifiers vs. F1 scores
Table 14: Correlation Coefficients (Pearson / Spearman) of $\mathbf{PAD}$ predictions from different classifiers vs. F1 scores.
Classifier Correlation Coefficients (P/S)
Logistic Regression .65 / .53
Random Forest .75 / .64
2-layer MLP .56 / .43
In summary, according to table 14, classifier choice plays a critical role in $\mathbf{PAD}$ prediction performance, with the random forest model demonstrating the highest correlation with F1 scores. This may be due to the ensemble nature of both the random forest and the sentiment analysis classifier (XGBoost), possibly enhancing alignment between predictions and F1 outcomes, while the MLP shows weaker correspondence.
B.8 Trade-off analysis between Lens and diversity measure
Figure 3: Pearson/Spearman correlation coefficients of the hybrid score vs. alpha. Both correlation coefficients are monotonically decreasing as alpha increases (more Lens impact and less $\mathbf{MDM}$ impact).
To further understand the interplay between Lens and the diversity measure, we conducted an experiment using a ”hybrid” score, defined as $\alpha\cdot\textsc{Lens}{}+(1-\alpha)\cdot\mathbf{MDM}$. As shown in fig. 3, our analysis reveals that increasing $\alpha$—thereby placing greater weight on LENS and less on diversity—results in a monotonic decrease in both the Pearson and Spearman correlation coefficients. This trend suggests that the LENS and diversity measures do not demonstrate synergistic effects when combined.
B.9 Ablation on the number of samples used for scoring on simple task
Table 15: Pearson / Spearman correlation between the Lens score and the test F1 score for different metrics with differet number of real instances used for scoring (100 and 200).
$|\mathcal{U}_{|}=100$ $|\mathcal{U}_{|}=200$
Lens $\mathbf{PAD}$ $\mathbf{MMD}^{2}$ $\mathbf{MDM}$ Mauve Lens $\mathbf{PAD}$ $\mathbf{MMD}^{2}$ $\mathbf{MDM}$ Mauve
-.03/-.13 .65/.47 .68/.44 .85/.70 .54/.51 .26/.38 .65/.53 .67/.45 .85/.68 .57/.53
We investigate the sensitivity of LENS proxies to the size of the unannotated real dataset $|\mathcal{U}_{r}|$ for the sentiment analysis task.As shown in table 15, representation-based metrics ($\mathbf{PAD}$, $\mathbf{MMD}^{2}$, $\mathbf{MDM}$, Mauve) demonstrate remarkable robustness in this domain, maintaining strong correlations (e.g., $\mathbf{MDM}$ Spearman $0.70$) even with as few as 100 samples. This aligns with our hypothesis that embedding spaces effectively capture distributional shifts in simpler, stylistic tasks. In contrast, Lens requires a larger sample size to stabilize for this specific task 5, exhibiting weak correlations at $|\mathcal{U}_{r}|=100$ (Pearson 0.03-0.03, Spearman 0.13-0.13) before recovering significantly to a positive Spearman correlation of $0.38$ at $|\mathcal{U}_{r}|=200$. This behavior highlights a distinct complexity trade-off: while representation metrics are highly data-efficient for stylistic domains, they fail on complex, long-horizon tasks such as Web Navigation. Conversely, Lens proves robust in those more complex domains even with very limited unlabled real data.